# Towards Intelligible Human-Robot Interaction: An Active Inference Approach to Occluded Pedestrian Scenarios

### v1

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 21st ACM/IEEE International Conference on Human-Robot Interaction; March 16–19, 2026; Edinburgh, Scotland, UKbooktitle: Proceedings of the 21st ACM/IEEE International Conference on Human-Robot Interaction (HRI ’26), March 16–19, 2026, Edinburgh, Scotland, UKdoi: 10.1145/3757279.3788670isbn: 979-8-4007-2128-1/2026/03
@@ 1. Introduction
Achieving safe interaction in complex driving scenarios remains a significant challenge, particularly in highly uncertain situations, such as occlusions. Although data-driven methods, including reinforcement learning, imitation learning, and large language models, have shown considerable success (Zhao et al., 2024; Huang et al., 2025; Zhou et al., 2024a, b), they often encounter substantial hurdles. These include the scarcity of data for safety-critical corner cases (Kiran et al., 2022; Chen et al., 2024) and a lack of interpretability (Atakishiyev et al., 2024). Consequently, several mechanism-inspired approaches (Engström et al., 2024a; Albrecht et al., 2021; Jiang and Wang, 2022) have been proposed. Among these, the free energy principle (Friston, 2010) and its process theory, active inference (Parr et al., 2022; Friston et al., 2017), stand out as theoretically rigorous and promising. They offer a biologically inspired framework with applications spanning robotics (Pio-Lopez et al., 2016), autonomous driving (Engström et al., 2024b; Schumann et al., 2025; Wei et al., 2025, 2024a), and human-computer interaction (HCI) (Stein et al., 2025). Recently, researchers (Lu et al., 2025) have characterized the exploratory applications of active inference in traffic occlusion scenarios (Engström et al., 2024b) as a form of mechanism-inspired human-like driving, positioning it as a paradigm for autonomous decision-making. Furthermore, a recent review (Murray-Smith et al., 2024) highlights active inference as a suitable framework for modeling the human-computer interaction loop.
@@ 3.3. Active Inference Interaction Framework
This is a comprehensive multi-part diagram illustrating an autonomous interaction framework, organized into three main sections: a top row of high-level flowcharts, a bottom row of concrete scenario illustrations, and a legend at the very bottom. The top section is divided into three interconnected parts. On the top-left, a light green box titled "Planning as Inference" shows a "CEM Iteration" loop. This loop is fed by a "Belief" and a highlighted "Hypothesis Injection" module, and it outputs a "Policy." In the top-center, a simple cycle of four ovals illustrates the core feedback loop: Action leads to Environment, which leads to Observation, which leads to Belief, and back to Action. On the top-right, a light yellow box titled "Belief Updating" shows how a "Belief" becomes a "Predicted Belief," which is corrected after conflicting with an "Observation." A key component in this correction phase is a highlighted "Conditional Reset" module, which feeds into a "Filter" to produce the new Belief. The bottom section contains two panels that provide visual examples. The panel on the bottom-left, titled "Conditional Reset," illustrates a scenario on a two-lane road where a truck occludes a pedestrian. It depicts a three-step process. Step 1 shows a belief about the pedestrian’s current position, labeled as "s-hat with subscript p-comma-zero and superscript n," being projected forward. Step 2 shows a conflict, marked with a red X, where the predicted future position is in an occluded area. Step 3 shows the belief being reset to its original position, preserving the hypothesis of the hidden pedestrian. The panel on the bottom-right, titled "Hypothesis Injection," shows a car planning a trajectory to avoid a pedestrian. An inset box details the hypotheses being considered: "Surge," "Reverse," and "Freeze," each with a corresponding pedestrian icon and motion arrow. At the very bottom, a legend defines the graphical elements: A red-filled circle is "Belief of current pedestrian position." An orange-outlined circle is "Belief of future pedestrian position." Blue bars are "Belief of current pedestrian velocity." Teal bars are "Belief of future pedestrian velocity." A thick green arrow is "Planned future trajectory."
@@ 4.2. Ablation Studies
0.0 0.8 $905\pm 0.012 0.0$$095\pm 0.012$ $\mathbf{7.974\pm 0.026}$ $3.679\pm 0.049$ $\mathbf{1.739\pm 0.041}$
@@
0.3 0.8 $947\pm 0.010 0.0$$053\pm 0.010$ $8.653\pm 0.052$ $\mathbf{3.890\pm 0.041}$ $1.540\pm 0.030$
0.8 0.8 $\mathbf{0.977\pm 0.007}$ $\mathbf{0.023\pm 0.007}$ $9.148\pm 0.064$ $3.367\pm 0.024$ $0.963\pm 0.016$
0.3 0.0 $0.250\pm 0.018$ $0.750\pm 0.018$ $\mathbf{6.448\pm 0.007}$ $3.770\pm 0.084$ $1.220\pm 0.045$
@@
0.3 0.8 $\mathbf{0.947\pm 0.010}$ $\mathbf{0.053\pm 0.010}$ $8.653\pm 0.052$ $\mathbf{3.890\pm 0.041}$ $\mathbf{1.540\pm 0.030}$
@@
A figure containing two line plots, labeled (a) and (b), comparing two conditions: "Without Conditional Reset" shown as a blue dashed line with circle markers, and "With Conditional Reset" shown as an orange dashed line with triangle markers.
Plot (a) shows "B(zp) versus Time (s)". The y-axis, B(zp), ranges from 0.00 to 1.00. The x-axis, Time, ranges from 0.0 to 10.0 seconds. The blue line starts near 0.4, decreases to 0 at 3 seconds, and then sharply jumps to 1.0 at approximately 4.2 seconds, where a vertical line indicates a "Collision". The orange line starts near 0.4 and remains constant until it also jumps to 1.0 at 4.2 seconds. A second vertical line near 9.5 seconds indicates "Goal Reached" for the orange trajectory.
Plot (b) shows "Speed (m/s) versus Time (s)". The y-axis, Speed, ranges from 0.0 to 12.5 m/s. The blue line starts around 10 m/s, dips slightly, then rises to a peak above 12.5 m/s at the 4.2-second "Collision" line, where the data ends. The orange line starts around 10.5 m/s and steadily decreases to a minimum of about 2.5 m/s at 4.2 seconds. After this point, the speed recovers and increases toward the "Goal Reached" event.
@@
As shown in Figure 3, the outcomes differ dramatically. The agent without Conditional Reset (blue curves) allows its belief in the pedestrian to decay to zero while the view is occluded (Figure 3a). This erroneous belief causes it to accelerate into the uncertain area (Figure 3b). When the pedestrian is finally observed at $t\approx 4.2\penalty 10000\ \mathrm{s}$, the agent is moving too fast to react, resulting in a collision. In contrast, the agent equipped with Conditional Reset (orange curves) maintains a persistent, non-zero belief about the potential hazard. This sustained caution compels it to decelerate as it approaches the occlusion, reaching a safe speed before the pedestrian is revealed. This cautious behavior allows it to easily avoid a collision and proceed to its goal, demonstrating that the mechanism is critical for safe interaction in uncertain environments.
@@
A figure containing two plots, labeled (a) and (b), sharing a common legend at the top. The legend shows four conditions: rho_H=0.0 (blue line with circle marker), rho_H=0.1 (light blue line with triangle marker), rho_H=0.3 (green line with square marker), and rho_H=0.8 (tan line with plus marker). Plot (a) is a top-down view of a driving scenario. The axes are "X Position (m)" from -5 to 15 and "Y Position (m)" from -10 to 5. An orange truck is depicted near the center. Four colored lines show vehicle trajectories starting from the left. The blue line proceeds almost straight and ends at a red ’X’ marker, indicating a collision with a red pedestrian icon emerging from behind the truck. The other lines (light blue, green, tan) swerve progressively more to the left, creating more space around the pedestrian. Plot (b) is a line graph of "Speed (m/s)" versus "Time (s)". The y-axis ranges from 0 to 12.5, and the x-axis from 0 to 8. All lines start around 10 m/s and show a dip in speed. The blue line dips shallowly before rising and abruptly ending at a vertical dash-dotted line labeled "Collision" around 4.8 seconds. The other three lines show a much deeper dip in speed, reaching a minimum of around 2.5 m/s, before recovering.
@@ 4.4.1. Baseline Methods
PPO-LSTM $0.725\pm 0.018$ $0.275\pm 0.018$ $\mathbf{4.188\pm 0.002}$ $\mathbf{4.080\pm 0.038}$ $\mathbf{9.663\pm 0.086}$
Ours $\mathbf{0.947\pm 0.010}$ $\mathbf{0.053\pm 0.010}$ $8.653\pm 0.052$ $3.890\pm 0.041$ $1.540\pm 0.030$
@@ 5.1. Pathway to Deployment
A potential deployment of our framework within a modular autonomous driving system would involve an upstream perception module that processes sensor data to detect objects like the occluding bus. Crucially, this is augmented by a sophisticated scene understanding model, such as a Vision-Language Model (VLM) (Pan et al., 2024; Zhou et al., 2024a), to provide high-level semantic context. This VLM’s role is to analyze the scene and infer the probability of latent hazards—for instance, assigning a high probability to an occluded pedestrian scenario upon seeing a bus stopped at a crosswalk. This output directly serves as the Initial Presence Belief ($B_{0}(z_{p})$) for our planner. Our belief-aware planning module then uses this structured, probabilistic world representation and computes a kinematically feasible trajectory that optimally balances safety and efficiency. Finally, this trajectory is passed to a downstream low-level controller for execution via precise steering and throttle commands. A key strength of this architecture is its two intuitive parameters for calibrating risk sensitivity. The VLM-generated $B_{0}(z_{p})$ allows the vehicle’s caution to dynamically adapt to real-time perceptual context. Concurrently, the Hypothesis Injection Ratio ($\rho_{H}$) acts as a configurable parameter for system designers to set the planner’s intrinsic defensiveness, tuning how proactively it plans for potential pedestrian movements. This dual-parameter system enables manufacturers to define distinct driving profiles to align the vehicle’s interactive behavior with specific safety requirements and user preferences.
@@ Appendix C Details of Ablation Studies
0.0 0.8 $0.783\pm 0.038$ $0.217\pm 0.038$ $\mathbf{7.685\pm 0.056}$ $\mathbf{3.806\pm 0.125}$ $\mathbf{1.798\pm 0.113}$
@@
0.3 0.8 $\mathbf{0.925\pm 0.024}$ $\mathbf{0.075\pm 0.024}$ $8.172\pm 0.076$ $3.651\pm 0.093$ $1.519\pm 0.081$
@@
0.3 0.0 $0.433\pm 0.045$ $0.567\pm 0.045$ $\mathbf{6.452\pm 0.013}$ $\mathbf{3.687\pm 0.140}$ $1.173\pm 0.080$
@@
0.3 0.8 $\mathbf{0.925\pm 0.024}$ $\mathbf{0.075\pm 0.024}$ $8.172\pm 0.076$ $3.651\pm 0.093$ $\mathbf{1.519\pm 0.081}$
@@
0.0 0.8 $0.808\pm 0.036$ $0.192\pm 0.036$ $\mathbf{7.721\pm 0.052}$ $3.723\pm 0.072$ $1.799\pm 0.074$
0.1 0.8 $0.758\pm 0.039$ $0.242\pm 0.039$ $7.796\pm 0.058$ $4.144\pm 0.083$ $\mathbf{1.838\pm 0.066}$
0.3 0.8 $0.850\pm 0.033$ $0.150\pm 0.033$ $8.692\pm 0.125$ $\mathbf{4.249\pm 0.093}$ $1.637\pm 0.065$
0.8 0.8 $\mathbf{0.983\pm 0.014}$ $\mathbf{0.017\pm 0.012}$ $10.431\pm 0.139$ $3.278\pm 0.040$ $0.872\pm 0.031$
0.3 0.0 $0.167\pm 0.034$ $0.833\pm 0.034$ $\mathbf{6.435\pm 0.018}$ $3.141\pm 0.103$ $0.954\pm 0.153$
@@
0.3 0.8 $\mathbf{0.850\pm 0.033}$ $\mathbf{0.150\pm 0.033}$ $8.692\pm 0.125$ $\mathbf{4.249\pm 0.093}$ $\mathbf{1.637\pm 0.065}$
@@
0.1 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $\mathbf{8.116\pm 0.038}$ $2.961\pm 0.030$ $\mathbf{1.212\pm 0.066}$
0.3 0.8 $992\pm 0.008 0.0$$008\pm 0.008$ $8.336\pm 0.051$ $\mathbf{3.267\pm 0.023}$ $1.056\pm 0.044$
0.8 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $8.363\pm 0.058$ $3.241\pm 0.017$ $0.828\pm 0.015$
0.3 0.0 $0.042\pm 0.018$ $0.958\pm 0.018$ $\mathbf{6.420\pm 0.033}$ $2.613\pm 0.143$ $0.490\pm 0.244$
@@
0.3 0.8 $\mathbf{0.992\pm 0.008}$ $\mathbf{0.008\pm 0.008}$ $8.336\pm 0.051$ $\mathbf{3.267\pm 0.023}$ $\mathbf{1.056\pm 0.044}$
@@
0.0 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $\mathbf{8.586\pm 0.033}$ $3.213\pm 0.042$ $\mathbf{1.636\pm 0.063}$
0.1 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $8.790\pm 0.054$ $3.399\pm 0.042$ $1.511\pm 0.063$
0.3 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $9.832\pm 0.129$ $\mathbf{3.661\pm 0.051}$ $1.475\pm 0.058$
@@
0.3 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $9.832\pm 0.129$ $\mathbf{3.661\pm 0.051}$ $\mathbf{1.475\pm 0.058}$
@@
0.0 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $\mathbf{7.537\pm 0.031}$ $4.913\pm 0.113$ $\mathbf{2.496\pm 0.033}$
0.1 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $7.610\pm 0.034$ $\mathbf{5.076\pm 0.111}$ $2.463\pm 0.033$
@@
0.8 0.8 $\mathbf{1.000\pm 0.000}$ $\mathbf{0.000\pm 0.000}$ $9.227\pm 0.143$ $3.568\pm 0.078$ $1.183\pm 0.037$
0.3 0.0 $0.608\pm 0.045$ $0.392\pm 0.045$ $\mathbf{6.451\pm 0.010}$ $4.080\pm 0.121$ $1.377\pm 0.046$
@@
0.3 0.4 $\mathbf{0.992\pm 0.008}$ $\mathbf{0.008\pm 0.008}$ $7.329\pm 0.044$ $4.352\pm 0.092$ $1.576\pm 0.046$
0.3 0.8 $967\pm 0.016 0.0$$033\pm 0.016$ $8.186\pm 0.094$ $\mathbf{4.678\pm 0.107}$ $\mathbf{2.039\pm 0.055}$

### v2

# arXiv:2602.23109v2[cs.RO] 27 Feb 2026

@@ Abstract.
journalyear: 2026conference: Proceedings of the 21st ACM/IEEE International Conference on Human-Robot Interaction; March 16–19, 2026; Edinburgh, Scotland, UKbooktitle: Proceedings of the 21st ACM/IEEE International Conference on Human-Robot Interaction (HRI ’26), March 16–19, 2026, Edinburgh, Scotland, UKdoi: 10.1145/3757279.3788670isbn: 979-8-4007-2128-1/2026/03
@@ 1. Introduction
Achieving safe interaction in complex driving scenarios remains a significant challenge, particularly in highly uncertain situations, such as occlusions. Although data-driven methods, including reinforcement learning, imitation learning, and large language models, have shown considerable success (Zhao et al., 2024; Huang et al., 2025; Zhou et al., 2024b, a), they often encounter substantial hurdles. These include the scarcity of data for safety-critical corner cases (Kiran et al., 2022; Chen et al., 2024) and a lack of interpretability (Atakishiyev et al., 2024). Consequently, several mechanism-inspired approaches (Engström et al., 2024a; Albrecht et al., 2021; Jiang and Wang, 2022) have been proposed. Among these, the free energy principle (Friston, 2010) and its process theory, active inference (Parr et al., 2022; Friston et al., 2017), stand out as theoretically rigorous and promising. They offer a biologically inspired framework with applications spanning robotics (Pio-Lopez et al., 2016), autonomous driving (Engström et al., 2024b; Schumann et al., 2025; Wei et al., 2025, 2024a), and human-computer interaction (HCI) (Stein et al., 2025). Recently, researchers (Lu et al., 2025) have characterized the exploratory applications of active inference in traffic occlusion scenarios (Engström et al., 2024b) as a form of mechanism-inspired human-like driving, positioning it as a paradigm for autonomous decision-making. Furthermore, a recent review (Murray-Smith et al., 2024) highlights active inference as a suitable framework for modeling the human-computer interaction loop.
@@ 3.3. Active Inference Interaction Framework
This is a comprehensive multi-part diagram illustrating an autonomous interaction framework, organized into three main sections: a top row of high-level flowcharts, a bottom row of concrete scenario illustrations, and a legend at the very bottom. The top section is divided into three interconnected parts. On the top-left, a light green box titled ”Planning as Inference” shows a ”CEM Iteration” loop. This loop is fed by a ”Belief” and a highlighted ”Hypothesis Injection” module, and it outputs a ”Policy.” In the top-center, a simple cycle of four ovals illustrates the core feedback loop: Action leads to Environment, which leads to Observation, which leads to Belief, and back to Action. On the top-right, a light yellow box titled ”Belief Updating” shows how a ”Belief” becomes a ”Predicted Belief,” which is corrected after conflicting with an ”Observation.” A key component in this correction phase is a highlighted ”Conditional Reset” module, which feeds into a ”Filter” to produce the new Belief. The bottom section contains two panels that provide visual examples. The panel on the bottom-left, titled ”Conditional Reset,” illustrates a scenario on a two-lane road where a truck occludes a pedestrian. It depicts a three-step process. Step 1 shows a belief about the pedestrian’s current position, labeled as ”s-hat with subscript p-comma-zero and superscript n,” being projected forward. Step 2 shows a conflict, marked with a red X, where the predicted future position is in an occluded area. Step 3 shows the belief being reset to its original position, preserving the hypothesis of the hidden pedestrian. The panel on the bottom-right, titled ”Hypothesis Injection,” shows a car planning a trajectory to avoid a pedestrian. An inset box details the hypotheses being considered: ”Surge,” ”Reverse,” and ”Freeze,” each with a corresponding pedestrian icon and motion arrow. At the very bottom, a legend defines the graphical elements: A red-filled circle is ”Belief of current pedestrian position.” An orange-outlined circle is ”Belief of future pedestrian position.” Blue bars are ”Belief of current pedestrian velocity.” Teal bars are ”Belief of future pedestrian velocity.” A thick green arrow is ”Planned future trajectory.”
@@ 4.2. Ablation Studies
0.0 0.8 $905\pm 0.012 0.0$$095\pm 0.012$ $\textbf{7.974}\pm\textbf{0.026}$ $3.679\pm 0.049$ $\textbf{1.739}\pm\textbf{0.041}$
@@
0.3 0.8 $947\pm 0.010 0.0$$053\pm 0.010$ $8.653\pm 0.052$ $\textbf{3.890}\pm\textbf{0.041}$ $1.540\pm 0.030$
0.8 0.8 $\textbf{0.977}\pm\textbf{0.007}$ $\textbf{0.023}\pm\textbf{0.007}$ $9.148\pm 0.064$ $3.367\pm 0.024$ $0.963\pm 0.016$
0.3 0.0 $0.250\pm 0.018$ $0.750\pm 0.018$ $\textbf{6.448}\pm\textbf{0.007}$ $3.770\pm 0.084$ $1.220\pm 0.045$
@@
0.3 0.8 $\textbf{0.947}\pm\textbf{0.010}$ $\textbf{0.053}\pm\textbf{0.010}$ $8.653\pm 0.052$ $\textbf{3.890}\pm\textbf{0.041}$ $\textbf{1.540}\pm\textbf{0.030}$
@@
A figure containing two line plots, labeled (a) and (b), comparing two conditions: ”Without Conditional Reset” shown as a blue dashed line with circle markers, and ”With Conditional Reset” shown as an orange dashed line with triangle markers.
Plot (a) shows ”B(zp) versus Time (s)”. The y-axis, B(zp), ranges from 0.00 to 1.00. The x-axis, Time, ranges from 0.0 to 10.0 seconds. The blue line starts near 0.4, decreases to 0 at 3 seconds, and then sharply jumps to 1.0 at approximately 4.2 seconds, where a vertical line indicates a ”Collision”. The orange line starts near 0.4 and remains constant until it also jumps to 1.0 at 4.2 seconds. A second vertical line near 9.5 seconds indicates ”Goal Reached” for the orange trajectory.
Plot (b) shows ”Speed (m/s) versus Time (s)”. The y-axis, Speed, ranges from 0.0 to 12.5 m/s. The blue line starts around 10 m/s, dips slightly, then rises to a peak above 12.5 m/s at the 4.2-second ”Collision” line, where the data ends. The orange line starts around 10.5 m/s and steadily decreases to a minimum of about 2.5 m/s at 4.2 seconds. After this point, the speed recovers and increases toward the ”Goal Reached” event.
@@
As shown in Figure 3, the outcomes differ dramatically. The agent without Conditional Reset (blue curves) allows its belief in the pedestrian to decay to zero while the view is occluded (Figure 3a). This erroneous belief causes it to accelerate into the uncertain area (Figure 3b). When the pedestrian is finally observed at $t\approx 4.2~\mathrm{s}$, the agent is moving too fast to react, resulting in a collision. In contrast, the agent equipped with Conditional Reset (orange curves) maintains a persistent, non-zero belief about the potential hazard. This sustained caution compels it to decelerate as it approaches the occlusion, reaching a safe speed before the pedestrian is revealed. This cautious behavior allows it to easily avoid a collision and proceed to its goal, demonstrating that the mechanism is critical for safe interaction in uncertain environments.
@@
A figure containing two plots, labeled (a) and (b), sharing a common legend at the top. The legend shows four conditions: rho_H=0.0 (blue line with circle marker), rho_H=0.1 (light blue line with triangle marker), rho_H=0.3 (green line with square marker), and rho_H=0.8 (tan line with plus marker). Plot (a) is a top-down view of a driving scenario. The axes are ”X Position (m)” from -5 to 15 and ”Y Position (m)” from -10 to 5. An orange truck is depicted near the center. Four colored lines show vehicle trajectories starting from the left. The blue line proceeds almost straight and ends at a red ’X’ marker, indicating a collision with a red pedestrian icon emerging from behind the truck. The other lines (light blue, green, tan) swerve progressively more to the left, creating more space around the pedestrian. Plot (b) is a line graph of ”Speed (m/s)” versus ”Time (s)”. The y-axis ranges from 0 to 12.5, and the x-axis from 0 to 8. All lines start around 10 m/s and show a dip in speed. The blue line dips shallowly before rising and abruptly ending at a vertical dash-dotted line labeled ”Collision” around 4.8 seconds. The other three lines show a much deeper dip in speed, reaching a minimum of around 2.5 m/s, before recovering.
@@ 4.4.1. Baseline Methods
PPO-LSTM $0.725\pm 0.018$ $0.275\pm 0.018$ $\textbf{4.188}\pm\textbf{0.002}$ $\textbf{4.080}\pm\textbf{0.038}$ $\textbf{9.663}\pm\textbf{0.086}$
Ours $\textbf{0.947}\pm\textbf{0.010}$ $\textbf{0.053}\pm\textbf{0.010}$ $8.653\pm 0.052$ $3.890\pm 0.041$ $1.540\pm 0.030$
@@ 5.1. Pathway to Deployment
A potential deployment of our framework within a modular autonomous driving system would involve an upstream perception module that processes sensor data to detect objects like the occluding bus. Crucially, this is augmented by a sophisticated scene understanding model, such as a Vision-Language Model (VLM) (Pan et al., 2024; Zhou et al., 2024b), to provide high-level semantic context. This VLM’s role is to analyze the scene and infer the probability of latent hazards—for instance, assigning a high probability to an occluded pedestrian scenario upon seeing a bus stopped at a crosswalk. This output directly serves as the Initial Presence Belief ($B_{0}(z_{p})$) for our planner. Our belief-aware planning module then uses this structured, probabilistic world representation and computes a kinematically feasible trajectory that optimally balances safety and efficiency. Finally, this trajectory is passed to a downstream low-level controller for execution via precise steering and throttle commands. A key strength of this architecture is its two intuitive parameters for calibrating risk sensitivity. The VLM-generated $B_{0}(z_{p})$ allows the vehicle’s caution to dynamically adapt to real-time perceptual context. Concurrently, the Hypothesis Injection Ratio ($\rho_{H}$) acts as a configurable parameter for system designers to set the planner’s intrinsic defensiveness, tuning how proactively it plans for potential pedestrian movements. This dual-parameter system enables manufacturers to define distinct driving profiles to align the vehicle’s interactive behavior with specific safety requirements and user preferences.
@@ Appendix C Details of Ablation Studies
0.0 0.8 $0.783\pm 0.038$ $0.217\pm 0.038$ $\textbf{7.685}\pm\textbf{0.056}$ $\textbf{3.806}\pm\textbf{0.125}$ $\textbf{1.798}\pm\textbf{0.113}$
@@
0.3 0.8 $\textbf{0.925}\pm\textbf{0.024}$ $\textbf{0.075}\pm\textbf{0.024}$ $8.172\pm 0.076$ $3.651\pm 0.093$ $1.519\pm 0.081$
@@
0.3 0.0 $0.433\pm 0.045$ $0.567\pm 0.045$ $\textbf{6.452}\pm\textbf{0.013}$ $\textbf{3.687}\pm\textbf{0.140}$ $1.173\pm 0.080$
@@
0.3 0.8 $\textbf{0.925}\pm\textbf{0.024}$ $\textbf{0.075}\pm\textbf{0.024}$ $8.172\pm 0.076$ $3.651\pm 0.093$ $\textbf{1.519}\pm\textbf{0.081}$
@@
0.0 0.8 $0.808\pm 0.036$ $0.192\pm 0.036$ $\textbf{7.721}\pm\textbf{0.052}$ $3.723\pm 0.072$ $1.799\pm 0.074$
0.1 0.8 $0.758\pm 0.039$ $0.242\pm 0.039$ $7.796\pm 0.058$ $4.144\pm 0.083$ $\textbf{1.838}\pm\textbf{0.066}$
0.3 0.8 $0.850\pm 0.033$ $0.150\pm 0.033$ $8.692\pm 0.125$ $\textbf{4.249}\pm\textbf{0.093}$ $1.637\pm 0.065$
0.8 0.8 $\textbf{0.983}\pm\textbf{0.014}$ $\textbf{0.017}\pm\textbf{0.012}$ $10.431\pm 0.139$ $3.278\pm 0.040$ $0.872\pm 0.031$
0.3 0.0 $0.167\pm 0.034$ $0.833\pm 0.034$ $\textbf{6.435}\pm\textbf{0.018}$ $3.141\pm 0.103$ $0.954\pm 0.153$
@@
0.3 0.8 $\textbf{0.850}\pm\textbf{0.033}$ $\textbf{0.150}\pm\textbf{0.033}$ $8.692\pm 0.125$ $\textbf{4.249}\pm\textbf{0.093}$ $\textbf{1.637}\pm\textbf{0.065}$
@@
0.1 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $\textbf{8.116}\pm\textbf{0.038}$ $2.961\pm 0.030$ $\textbf{1.212}\pm\textbf{0.066}$
0.3 0.8 $992\pm 0.008 0.0$$008\pm 0.008$ $8.336\pm 0.051$ $\textbf{3.267}\pm\textbf{0.023}$ $1.056\pm 0.044$
0.8 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $8.363\pm 0.058$ $3.241\pm 0.017$ $0.828\pm 0.015$
0.3 0.0 $0.042\pm 0.018$ $0.958\pm 0.018$ $\textbf{6.420}\pm\textbf{0.033}$ $2.613\pm 0.143$ $0.490\pm 0.244$
@@
0.3 0.8 $\textbf{0.992}\pm\textbf{0.008}$ $\textbf{0.008}\pm\textbf{0.008}$ $8.336\pm 0.051$ $\textbf{3.267}\pm\textbf{0.023}$ $\textbf{1.056}\pm\textbf{0.044}$
@@
0.0 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $\textbf{8.586}\pm\textbf{0.033}$ $3.213\pm 0.042$ $\textbf{1.636}\pm\textbf{0.063}$
0.1 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $8.790\pm 0.054$ $3.399\pm 0.042$ $1.511\pm 0.063$
0.3 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $9.832\pm 0.129$ $\textbf{3.661}\pm\textbf{0.051}$ $1.475\pm 0.058$
@@
0.3 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $9.832\pm 0.129$ $\textbf{3.661}\pm\textbf{0.051}$ $\textbf{1.475}\pm\textbf{0.058}$
@@
0.0 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $\textbf{7.537}\pm\textbf{0.031}$ $4.913\pm 0.113$ $\textbf{2.496}\pm\textbf{0.033}$
0.1 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $7.610\pm 0.034$ $\textbf{5.076}\pm\textbf{0.111}$ $2.463\pm 0.033$
@@
0.8 0.8 $\textbf{1.000}\pm\textbf{0.000}$ $\textbf{0.000}\pm\textbf{0.000}$ $9.227\pm 0.143$ $3.568\pm 0.078$ $1.183\pm 0.037$
0.3 0.0 $0.608\pm 0.045$ $0.392\pm 0.045$ $\textbf{6.451}\pm\textbf{0.010}$ $4.080\pm 0.121$ $1.377\pm 0.046$
@@
0.3 0.4 $\textbf{0.992}\pm\textbf{0.008}$ $\textbf{0.008}\pm\textbf{0.008}$ $7.329\pm 0.044$ $4.352\pm 0.092$ $1.576\pm 0.046$
0.3 0.8 $967\pm 0.016 0.0$$033\pm 0.016$ $8.186\pm 0.094$ $\textbf{4.678}\pm\textbf{0.107}$ $\textbf{2.039}\pm\textbf{0.055}$